OpenStudy (anonymous):

Evaluate each expression for the given value of the variable. 8p-12 ; p=4

OpenStudy (anonymous):

20

OpenStudy (anonymous):

substitue!!!

OpenStudy (paxpolaris):

8*4 - 12 = 32 - 12

OpenStudy (rulnick):

8p-12 at p=4 is 8(4)-12 = 32 -12 = 20.

OpenStudy (rulnick):

matt+krypton+pax=3*correct.